OpenStudy (anonymous):

2. Simplify and solve this equation: 4m + 9 + 5m – 12 = 42. A. m = -4 B. m = –5 C. m = 4 D. m = 5

OpenStudy (anonymous):

u have to do the equation

OpenStudy (anonymous):

first collect like terms, then isolate the variable

jimthompson5910 (jim_thompson5910):

4m+9+5m-12=42 9m-3=42 9m=42+3 9m=45 m=45/9 m=5

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw, glad to be of help